Identify the type of the differential equation y =xy'+1-lny' and hence
solve it
1
Expert's answer
2020-03-27T14:27:38-0400

"y=-ln(dy\/dx)+xdy\/dx+1"

This is Clairaut equation

Differentiate both sides with respect to x

"{\frac {dy} {dx}}=x{\frac {d^2y} {dx^2}}+{\frac {dy} {dx}}-{\frac {{\frac {d^2y} {dx^2}}} {{\frac {dy} {dx}}}}"

Collect in terms of "{\frac {d^2y} {dx^2}}"

"{\frac {dy} {dx}}={\frac {dy} {dx}}+{\frac {d^2y} {dx^2}}(x-{\frac {1} {{\frac {dy} {dx}}}})"

Then

"{\frac {d^2y} {dx^2}}(x-{\frac {1} {{\frac {dy} {dx}}}})=0"

Solve

"{\frac {d^2y} {dx^2}}=0" and "(x-{\frac {1} {{\frac {dy} {dx}}}})=0" separately:

For "{\frac {d^2y} {dx^2}}=0" :

"{\frac {dy} {dx}}=\int 0dx=C_1"

Substitute "y'=C_1" into initial equation

"y=-ln(C_1)+C_1x+1"

For "(x-{\frac {1} {{\frac {dy} {dx}}}})=0" :

"{\frac {dy} {dx}}={\frac {1} {x}}"

Substitute "y'=1\/x" into initial equation

"y=-ln({\frac {1} {x}})+2"


Answer: "y=-ln(C_1)+C_1x+1" or "y=-ln({\frac {1} {x}})+2"